OpenStudy (anonymous):

wat is 29/25 as a decimal

OpenStudy (anonymous):

25/25 = 1, leaves you 4/25

OpenStudy (anonymous):

do you have any answer choices?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

multiply 4/25 by 4/4 and it will be pretty easy to see the decimal.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

\[\large \frac{ 29 }{ 25 }=\frac{ 25 }{ 25 }+\frac{ 4 }{ 25 }=1+\frac{ 4 }{ 25 }*\frac{ 4 }{ 4 }=1.something\]

OpenStudy (anonymous):

yw